OpenStudy (anonymous):

this is a fill in the missing number -11+-28+-46+_____+20=-94 I took what was given and did the calculation then I subtracted -94 to get my missing answer. When I checked my answer it was incorrect. I got -65 as the missing number. Please advise. Thanks

OpenStudy (anonymous):

whoops hold on, messed up my bad -11-28 = -39 - 46 = -85 + 20 = -65 -94 + 65 =29

OpenStudy (angela210793):

let's say x is the missing number so u have -11+(-28)+(-46)+x+20=-94 -11-28-46+x+20=-94 x=-94+11+28+46-20 x=-29

OpenStudy (anonymous):

yeah there you go...

OpenStudy (anonymous):

ok but why did the signs on the number change when solving for x? example -11 to 11 -28 to 28 etc...

OpenStudy (angela210793):

u see when u have for example x-5=8 u have to add 5 to both sides so that u can get x(cause -a+a=0) x-5+5=8+5 x=13 so tht's what we did,we added 11,28,46 and -20 to both sides so tht we can have only x in the left side

OpenStudy (anonymous):

Oh yeah, I forgot that part....Thanks :-)

OpenStudy (angela210793):

ur welcome :)

OpenStudy (anonymous):

-12+-43+---- =-11 well i added up -12+-43 and got -55 then i took -55minus -11 and got 44 but im not sure if thats is the answer to the problem